Thursday, April 21, 2011

Nightlog 2011-04-20

SA: Alexei
SO: Siphelo
Others: Eric, Charl, Lisa


=> No observations tonight because major tracker failure.
Anyway weather was bad with strong wing, heavy clouds and high humidity

=> There were problems with SALTICAM, that we were not able to change filters.
The problem was fixed by Charl

=> PCON program gives very unstable result. We have got problems with
setup of all subsystems like changing grating, filters, masks,
grating rotation etc. We restarted program and it helps, but we still
have many problems.
In case of real observations our efficiency will be VERY LOW for this reason.

Very often we have "unknown light source problem", when, after changing
configuration, we do not see reference spectrum,
but just bright source in the right-up corner. Next exposure with the same
configuration produces "normal" result. See, for example, pairs of images:
P27 (bad) and P28 (good); P29 (bad) and P30 (good);
P32 (bad) and P33 (good).


No comments:

Post a Comment